OpenStudy (anonymous):

wILL GIVE MEDAL DONT GIVE ME THE ANSWER Which expression is equivalent to the expression Four and one-eighth divided by five and two-fifths? thirty-three eighths times twenty-seven fifths thirty-three eighths times five twenty-sevenths eight forty-firsts times fifty-two fifths forty-one eighths times five fifty-seconds

OpenStudy (anonymous):

Four and one-eighth = ? look at this : 2 and nine - three = 2*3+9/3 = 15/3=5 ok ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So what is the Four and one-eighth ? ( look at this : )  look at this : 32/3 = 10 and 2 - 3therd ok ?! got it ? :D

OpenStudy (anonymous):

I see

OpenStudy (anonymous):

so what is the : Four and one-eighth=?

OpenStudy (nerdgamer_):

@CrazyCountryGirl he did not say answer